# GRAND: Graph Release with Assured Node Differential Privacy

### v2

@@ B.2 Latent distribution consistency: one dimensional case
Our goal is to show that $\tilde{Z}_{i}\xrightarrow{d}Y\sim F$. We achieve this by leveraging Lemma B.3, which states that $\bar{Z}_{i}\xrightarrow{d}Y\sim F$. The remaining task is to prove that $\tilde{Z}_{i}-\bar{Z}_{i}\xrightarrow{P}0$. Specifically, we want to show that for any $\varepsilon>0$, $P\left(|\tilde{Z}_{i}-\bar{Z}_{i}|\geq\varepsilon\right)\xrightarrow{m,n\to\infty}0$.
@@
$u_{i}=G\left(\hat{F}_{m}(\hat{Z}_{i})+e_{i}\right),\quad v_{i}=G\left(\hat{F}_{m}(Z_{i})+e_{i}\right).$
Then, $|\tilde{Z}_{i}-\bar{Z}_{i}|=|\hat{F}_{m}^{-1}(u_{i})-\hat{F}_{m}^{-1}(v_{i})|$. Our strategy is to show that $|u_{i}-v_{i}|$ is small in probability, and then use the uniform equicontinuity of $\hat{F}_{m}^{-1}$ from Theorem B.1.
We first analyze $|u_{i}-v_{i}|=\left|G\left(\hat{F}_{m}(\hat{Z}_{i})+e_{i}\right)-G\left(\hat{F}_{m}(Z_{i})+e_{i}\right)\right|$. Since $G$ is a CDF of a continuous distribution (specifically, $U(0,1)+\text{Laplace}$), it is uniformly continuous on $\bf R$}. For any $\gamma>0$, there exists $\delta_{G,\gamma}>0$ such that for any $x,y\in{\bf R}$:
@@
Now, let’s bound the difference in the arguments of $G$: $|(\hat{F}_{m}(\hat{Z}_{i})+e_{i})-(\hat{F}_{m}(Z_{i})+e_{i})|=|\hat{F}_{m}(\hat{Z}_{i})-\hat{F}_{m}(Z_{i})|$. Since the kernel $K$ is bounded and Lipschitz continuous, $\hat{F}_{m}$ itself is Lipschitz continuous. Let $L_{\hat{F}_{m}}$ be its Lipschitz constant (proportional to $K_{\max}/h$). Then, on the event where $\sup_{1\leq j\leq n}|\hat{Z}_{j}-Z_{j}|\leq\delta_{m}$ (which holds with probability tending to 1 by assumption), we have:
$|\hat{F}_{m}(\hat{Z}_{i})-\hat{F}_{m}(Z_{i})|\leq L_{\hat{F}_{m}}|\hat{Z}_{i}-Z_{i}|\leq L_{\hat{F}_{m}}\delta_{m}.$
As $\delta_{m}\to 0$, it follows that $|\hat{F}_{m}(\hat{Z}_{i})-\hat{F}_{m}(Z_{i})|\xrightarrow{P}0$. This means that for any $\delta_{G,\gamma}>0$,
$P\left(|\hat{F}_{m}(\hat{Z}_{i})-\hat{F}_{m}(Z_{i})|\geq\delta_{G,\gamma}\right)\xrightarrow{m\to\infty}0.$
By (17), this implies $|u_{i}-v_{i}|\xrightarrow{P}0$.
Next, we handle the mapping of $\hat{F}_{m}^{-1}$.
From Theorem B.1, the sequence $\{\hat{F}_{m}^{-1}\}$ is uniformly equicontinuous in probability on any closed interval $a,b]\subset(0,1)$. Specifically, for any $\varepsilon>0$ and $\eta>0$, there exist $\delta^{\prime}>0$ and $M_{1}\in\mathbb{N}$ such that for all $m\geq M_{1}$:
$P\left(\sup_{\begin{subarray}{c}q_{1},q_{2}\in[a,b]\\ |q_{1}-q_{2}|<\delta^{\prime}\end{subarray}}|\hat{F}_{m}^{-1}(q_{1})-\hat{F}_{m}^{-1}(q_{2})|\geq\varepsilon\right)<\eta/3.$ (18)
The arguments to $\hat{F}_{m}^{-1}$ are $u_{i}=G(\hat{F}_{m}(\hat{Z}_{i})+e_{i})$ and $v_{i}=G(\hat{F}_{m}(Z_{i})+e_{i})$. Both $u_{i}$ and $v_{i}$ are random variables in $[0,1]$. We know $v_{i}\xrightarrow{d}U(0,1)$. Therefore, for any $\eta_{0}>0$, we can choose a closed interval $a,b]\subset(0,1)$ such that $P(v_{i}\in[a,b])\geq 1-\eta_{0}/2$. Since $|u_{i}-v_{i}|\xrightarrow{P}0$ (from Step 1), it also follows that $u_{i}\xrightarrow{P}v_{i}$. Consequently, $P(u_{i}\in[a,b])\geq 1-\eta_{0}/2$ for sufficiently large $m$. Thus, for any $\eta_{0}>0$, there exists an interval $[a,b]$ such that for sufficiently large $m$,
$P(u_{i}\in[a,b]\text{ and }v_{i}\in[a,b])\geq 1-\eta_{0}.$ (19)
@@
    Since $|\hat{F}_{m}(\hat{Z}_{i})-\hat{F}_{m}(Z_{i})|\xrightarrow{P}0$, there exists $M_{2}\in\mathbb{N}$ such that for $m\geq M_{2}$,
    $P\left(|\hat{F}_{m}(\hat{Z}_{i})-\hat{F}_{m}(Z_{i})|\geq\delta_{G,\gamma}\right)<\eta/3.$
@@
    Choose $\eta_{0}=\eta/3$ for (19), which defines the interval $[a,b]$ and ensures $P(u_{i}\in[a,b]\text{ and }v_{i}\in[a,b])\geq 1-\eta/3$ for $m\geq M_{3}$.
@@
$\mathscr{E}_{m}=\left\{|\hat{F}_{m}(\hat{Z}_{i})-\hat{F}_{m}(Z_{i})|<\delta_{G,\gamma}\right\}\cap\left\{u_{i}\in[a,b]\text{ and }v_{i}\in[a,b]\right\}\cap\left\{\sup_{\begin{subarray}{c}q_{1},q_{2}\in[a,b]\\ |q_{1}-q_{2}|<\delta^{\prime}\end{subarray}}|\hat{F}_{m}^{-1}(q_{1})-\hat{F}_{m}^{-1}(q_{2})|<\varepsilon\right\}.$
For $m\geq\max(M_{1},M_{2},M_{3})$, we have $P(\mathscr{E}_{m}^{c})\leq\eta/3+\eta/3+\eta/3=\eta$. On the event $\mathscr{E}_{m}$:
@@
$|\tilde{Z}_{i}-\bar{Z}_{i}|=|\hat{F}_{m}^{-1}(u_{i})-\hat{F}_{m}^{-1}(v_{i})|<\varepsilon.$
Since $P(\mathscr{E}_{m}^{c})\xrightarrow{m\to\infty}0$, we conclude that $|\tilde{Z}_{i}-\bar{Z}_{i}|\xrightarrow{P}0$.
@@ B.3 Latent distribution consistency: multidimensional case
Let $Z_{i}=(Z_{i1},Z_{i2})$ be i.i.d. $2$-dimensional random vectors from a distribution $F$ on ${\bf R}^{2}$. Assume that the true distribution $F$ has compact support $S\subset{\bf R}^{2}$. Assume the following conditions hold:
@@
  • The perturbation size $\delta_{m}$ satisfies $\delta_{m}=o((\log m)^{-c})$, and $\sup_{1\leq j\leq n}\|\hat{Z}_{j}-Z_{j}\|\leq\delta_{m}$ with probability tending to 1 as $m\to\infty$, for some constant $c>0$.
  • For each $k=1,2$, the bandwidth $h_{m}^{(k)}$ for $\hat{F}_{k|1:(k-1),m}$ is chosen as $h_{m}=h_{m}^{(k)}=(\log m)^{-c}$ for the same constant $c>0$.
@@
For notational simplicity within this proof, let $F_{k}(\cdot)$ denote the marginal CDF of $Z_{k}$, and $F_{k|1:(k-1)}(\cdot|u)$ denote the true conditional CDF for the $k$-th dimension given $u=(z_{1},\dots,z_{k-1})$. Similarly, $\hat{F}_{k,m}$ and $\hat{F}_{k|1:(k-1),m}(\cdot|u)$ are their estimators. Inverse functions are denoted with 1-1. Let $S_{k}$ represent the compact support of $Z_{k}$ (derived from the compact support $S$ of $F$), and $S_{1:(k-1)}$ be the compact support for the conditioning variables $(Z_{1},\dots,Z_{k-1})$. Because $F$ has compact support $S\subset{\bf R}^{d}$, all $Z_{ik}$ and $(Z_{i1},\dots,Z_{i,k-1})$ almost surely lie within their compact projected supports. This is crucial for applying uniform convergence results from previous lemmas, which hold over compact domains.
@@
$\tilde{Z}_{i1}=\hat{F}_{1,m}^{-1}\left(G\left(\hat{F}_{1,m}(\hat{Z}_{i1})+e_{i}^{(1)}\right)\right)$
@@
$\tilde{Z}_{i1}\xrightarrow{d}Y_{1}\sim F_{1}\quad\text{as }m\to\infty$
Step 2 – convergence of the second coordinate $\tilde{Z}_{i12}$ (conditioning on $\tilde{Z}_{i1}$):
We now analyze the conditional distribution of $\tilde{Z}_{i2}$ given $\tilde{Z}_{i1}=u_{1}$, for any fixed $u_{1}\in S_{1}$ (the compact support of $Y_{1}$). The quantity of interest is $P(\tilde{Z}_{i2}\leq x_{2}|\tilde{Z}_{i1}=u_{1})$. For this analysis, we treat $u_{1}$ as a deterministic conditioning variable. The definition of $\tilde{Z}_{i2}$ involves conditioning on $\tilde{Z}_{i1}$, so we consider the variable:
$\tilde{Z}_{i2}(u_{1})=\hat{F}_{2|1,m}^{-1}\left(G\left(\hat{F}_{2|1,m}(\hat{Z}_{i2}|\hat{Z}_{i1})+e_{i}^{(2)}\right)\right|u_{1}\Big{)}$
@@
$\bar{Z}_{i2}(u_{1})=\hat{F}_{2|1,m}^{-1}\left(G\left(\hat{F}_{2|1,m}(Z_{i2}|Z_{i1})+e_{i}^{(2)}\right)\right|u_{1}\Big{)}$
$T_{i2}(u_{1})=F_{2|1}^{-1}\left(G\left(F_{2|1}(Z_{i2}|Z_{i1})+e_{i}^{(2)}\right)\right|u_{1}\Big{)}$
Our goal is to show $\tilde{Z}_{i2}(u_{1})-\bar{Z}_{i2}(u_{1})\xrightarrow{P}0$ and $\bar{Z}_{i2}(u_{1})-T_{i2}(u_{1})\xrightarrow{P}0$.
To show $\bar{Z}_{i2}(u_{1})-T_{i2}(u_{1})\xrightarrow{P}0$, we directly apply the logic from Lemma B.4 to the conditional setting. For a fixed $u_{1}$, the functions $\hat{F}_{2|1,m}(\cdot|u_{1})$ and $F_{2|1}(\cdot|u_{1})$ are well-defined. Let $q_{C}=G(\hat{F}_{2|1,m}(Z_{i2}|Z_{i1})+e_{i}^{(2)})$ and $q_{D}=G(F_{2|1}(Z_{i2}|Z_{i1})+e_{i}^{(2)})$. We want to show $|\hat{F}_{2|1,m}^{-1}(q_{C}|u_{1})-F_{2|1}^{-1}(q_{D}|u_{1})|\xrightarrow{P}0$. This splits into:
$|\hat{F}_{2|1,m}^{-1}(q_{C}|u_{1})-F_{2|1}^{-1}(q_{C}|u_{1})|+|F_{2|1}^{-1}(q_{C}|u_{1})-F_{2|1}^{-1}(q_{D}|u_{1})|.$
  • Term 1: By Theorem B.2, $\sup_{u\in S_{1},q\in[a,b]}|\hat{F}_{2|1,m}^{-1}(q|u)-F_{2|1}^{-1}(q|u)|\xrightarrow{P}0$. Since $u_{1}$ is fixed in $S_{1}$ and $q_{C}$ is a random variable in $[0,1]$, this term goes to 0 in probability, using the same type of proof as in Theorem B.2.
  • Term 2: By uniform continuity of $F_{2|1}^{-1}(\cdot|u_{1})$ w.r.t. its first argument, this term goes to 0 in probability if $|q_{C}-q_{D}|\xrightarrow{P}0$. $|q_{C}-q_{D}|=|G(\hat{F}_{2|1,m}(Z_{i2}|Z_{i1})+e_{i}^{(2)})-G(F_{2|1}(Z_{i2}|Z_{i1})+e_{i}^{(2)})|$. By uniform continuity of $G$, this goes to 0 if $|\hat{F}_{2|1,m}(Z_{i2}|Z_{i1})-F_{2|1}(Z_{i2}|Z_{i1})|\xrightarrow{P}0$. This is true by Lemma B.2.
Combining the previous two results, we have $\bar{Z}_{i2}(u_{1})-T_{i2}(u_{1})\xrightarrow{P}0$.
On the other hand, the term $\tilde{Z}_{i2}(u_{1})-\bar{Z}_{i2}(u_{1})\xrightarrow{P}0$ involves changes in the inner argument’s conditioning variable ($\hat{Z}_{i1}$ to $Z_{i1}$) and the value being fed into $G$ ($\hat{Z}_{i2}$ to $Z_{i2}$).
$|\tilde{Z}_{i2}(u_{1})-\bar{Z}_{i2}(u_{1})|=\left|\hat{F}_{2|1,m}^{-1}\left(G\left(\hat{F}_{2|1,m}(\hat{Z}_{i2}|\hat{Z}_{i1})+e_{i}^{(2)}\right)\right|u_{1}\Big{)}-\hat{F}_{2|1,m}^{-1}\left(G\left(\hat{F}_{2|1,m}(Z_{i2}|Z_{i1})+e_{i}^{(2)}\right)\right|u_{1}\Big{)}\right|$
@@
$|\hat{F}_{2|1,m}^{-1}(q_{A}|u_{1})-\hat{F}_{2|1,m}^{-1}(q_{B}|u_{1})|.$
By the uniform equicontinuity of $\hat{F}_{2|1,m}^{-1}(\cdot|u_{1})$ with respect to its first argument from Theorem B.2, this term will converge to 0 in probability if $|q_{A}-q_{B}|\xrightarrow{P}0$.
@@
$|q_{A}-q_{B}|=\left|G\left(\hat{F}_{2|1,m}(\hat{Z}_{i2}|\hat{Z}_{i1})+e_{i}^{(2)}\right)-G\left(\hat{F}_{2|1,m}(Z_{i2}|Z_{i1})+e_{i}^{(2)}\right)\right|$
By uniform continuity of $G$, it is sufficient to show $|\hat{F}_{2|1,m}(\hat{Z}_{i2}|\hat{Z}_{i1})-\hat{F}_{2|1,m}(Z_{i2}|Z_{i1})|\xrightarrow{P}0$. By plugging in an intermediate term and calling the triangle inequality, we have
$|\hat{F}_{2|1,m}(\hat{Z}_{i2}|\hat{Z}_{i1})-\hat{F}_{2|1,m}(Z_{i2}|Z_{i1})|\leq|\hat{F}_{2|1,m}(\hat{Z}_{i2}|\hat{Z}_{i1})-\hat{F}_{2|1,m}(Z_{i2}|\hat{Z}_{i1})|+|\hat{F}_{2|1,m}(Z_{i2}|\hat{Z}_{i1})-\hat{F}_{2|1,m}(Z_{i2}|Z_{i1})|$

### v3

@@ B.2 Latent distribution consistency: one dimensional case
Our goal is to show that $\tilde{Z}_{i}\xrightarrow{d}Y\sim F$. We achieve this by leveraging Lemma B.3, which states that $\bar{Z}_{i}\xrightarrow{d}Y\sim F$. The remaining task is to prove that $\tilde{Z}_{i}-\bar{Z}_{i}\xrightarrow{\mathbb{P}}0$. Specifically, we want to show that for any $\varepsilon>0$, $\mathbb{P}(|\tilde{Z}_{i}-\bar{Z}_{i}|\geq\varepsilon)\xrightarrow{m\to\infty}0$.
@@
$u_{i}=G(\hat{F}_{m}(\hat{Z}_{i})+e_{i}),\quad v_{i}=G(\hat{F}_{m}(Z_{i})+e_{i}).$
Then, $|\tilde{Z}_{i}-\bar{Z}_{i}|=\big|\hat{F}_{m}^{-1}(u_{i})-\hat{F}_{m}^{-1}(v_{i})\big|$. Our strategy is to show that $|u_{i}-v_{i}|$ is small in probability, and then use the uniform equicontinuity of $\hat{F}_{m}^{-1}$ from Theorem B.1.
We first analyze $|u_{i}-v_{i}|=\big|G(\hat{F}_{m}(\hat{Z}_{i})+e_{i})-G(\hat{F}_{m}(Z_{i})+e_{i})\big|$. Since $G$ is a CDF of a continuous distribution (specifically, $\text{Uniform}(0,1)+\text{Laplace}(0,1/\varepsilon)$), it is uniformly continuous on $\mathbb{R}$. For any $\gamma>0$, there exists $\delta_{G,\gamma}>0$ such that for any $x,y\in\mathbb{R}$:
@@
Now, let’s bound the difference in the arguments of $G$: $\big|(\hat{F}_{m}(\hat{Z}_{i})+e_{i})-(\hat{F}_{m}(Z_{i})+e_{i})\big|=\big|\hat{F}_{m}(\hat{Z}_{i})-\hat{F}_{m}(Z_{i})\big|$. Since the kernel $K$ is bounded and Lipschitz continuous, $\hat{F}_{m}$ itself is Lipschitz continuous. Let $L_{\hat{F}_{m}}$ be its Lipschitz constant (proportional to $K_{\max}/h$). Then, on the event where $\sup_{1\leq i\leq n}|\hat{Z}_{i}-Z_{i}|\leq\delta_{m}$ (which holds with probability tending to 1 by assumption), we have:
$\big|\hat{F}_{m}(\hat{Z}_{i})-\hat{F}_{m}(Z_{i})\big|\leq L_{\hat{F}_{m}}|\hat{Z}_{i}-Z_{i}|\leq L_{\hat{F}_{m}}\delta_{m}.$
As $\delta_{m}\to 0$, it follows that $\big|\hat{F}_{m}(\hat{Z}_{i})-\hat{F}_{m}(Z_{i})\big|\xrightarrow{\mathbb{P}}0$. This means that for any $\delta_{G,\gamma}>0$,
$\mathbb{P}\left(\big|\hat{F}_{m}(\hat{Z}_{i})-\hat{F}_{m}(Z_{i})\big|\geq\delta_{G,\gamma}\right)\xrightarrow{m\to\infty}0.$
By (17), this implies $|u_{i}-v_{i}|\xrightarrow{\mathbb{P}}0$.
Next, we handle the mapping of $\hat{F}_{m}^{-1}$. From Theorem B.1, the sequence $\{\hat{F}_{m}^{-1}\}$ is uniformly equicontinuous in probability on any closed interval $a,b]\subset(0,1)$. Specifically, for any $\varepsilon>0$ and $\eta>0$, there exist $\delta^{\prime}>0$ and $M_{1}\in\mathbb{N}$ such that for all $m\geq M_{1}$:
$\mathbb{P}\left(\sup_{\begin{subarray}{c}q_{1},q_{2}\in[a,b]\\ |q_{1}-q_{2}|<\delta^{\prime}\end{subarray}}\big|\hat{F}_{m}^{-1}(q_{1})-\hat{F}_{m}^{-1}(q_{2})\big|\geq\varepsilon\right)<\frac{\eta}{3}.$ (18)
The arguments to $\hat{F}_{m}^{-1}$ are $u_{i}=G(\hat{F}_{m}(\hat{Z}_{i})+e_{i})$ and $v_{i}=G(\hat{F}_{m}(Z_{i})+e_{i})$. Both $u_{i}$ and $v_{i}$ are random variables in $[0,1]$. We know $v_{i}\xrightarrow{d}\text{Uniform}(0,1)$. Therefore, for any $\eta_{0}>0$, we can choose a closed interval $a,b]\subset(0,1)$ such that $\mathbb{P}(v_{i}\in[a,b])\geq 1-\eta_{0}/2$. Since $|u_{i}-v_{i}|\xrightarrow{\mathbb{P}}0$ (from Step 1), it also follows that $u_{i}\xrightarrow{\mathbb{P}}v_{i}$. Consequently, $P(u_{i}\in[a,b])\geq 1-\eta_{0}/2$ for sufficiently large $m$. Thus, for any $\eta_{0}>0$, there exists an interval $[a,b]$ such that for sufficiently large $m$,
$\mathbb{P}(u_{i}\in[a,b]\text{ and }v_{i}\in[a,b])\geq 1-\eta_{0}.$ (19)
@@
    Since $\big|\hat{F}_{m}(\hat{Z}_{i})-\hat{F}_{m}(Z_{i})\big|\xrightarrow{\mathbb{P}}0$, there exists $M_{2}\in\mathbb{N}$ such that for $m\geq M_{2}$,
    $\mathbb{P}\left(\big|\hat{F}_{m}(\hat{Z}_{i})-\hat{F}_{m}(Z_{i})\big|\geq\delta_{G,\gamma}\right)<\frac{\eta}{3}.$
@@
    Choose $\eta_{0}=\eta/3$ for (19), which defines the interval $[a,b]$ and ensures $\mathbb{P}(u_{i}\in[a,b]\text{ and }v_{i}\in[a,b])\geq 1-\eta/3$ for $m\geq M_{3}$.
@@
$\mathscr{E}_{m}=\left\{\big|\hat{F}_{m}(\hat{Z}_{i})-\hat{F}_{m}(Z_{i})\big|<\delta_{G,\gamma}\right\}\cap\left\{u_{i}\in[a,b]\text{ and }v_{i}\in[a,b]\right\}\cap\left\{\sup_{\begin{subarray}{c}q_{1},q_{2}\in[a,b]\\ |q_{1}-q_{2}|<\delta^{\prime}\end{subarray}}\big|\hat{F}_{m}^{-1}(q_{1})-\hat{F}_{m}^{-1}(q_{2})\big|<\varepsilon\right\}.$
For $m\geq\max\{M_{1},M_{2},M_{3}\}$, we have $\mathbb{P}(\mathscr{E}_{m}^{c})\leq\eta/3+\eta/3+\eta/3=\eta$. On the event $\mathscr{E}_{m}$:
@@
$|\tilde{Z}_{i}-\bar{Z}_{i}|=\big|\hat{F}_{m}^{-1}(u_{i})-\hat{F}_{m}^{-1}(v_{i})\big|<\varepsilon.$
Since $\mathbb{P}(\mathscr{E}_{m}^{c})\xrightarrow{m\to\infty}0$, we conclude that $|\tilde{Z}_{i}-\bar{Z}_{i}|\xrightarrow{\mathbb{P}}0$.
@@ B.3 Latent distribution consistency: multidimensional case
Let $Z_{i}=(Z_{i1},Z_{i2})$ be i.i.d. $2$-dimensional random vectors from a distribution $F$ on $\mathbb{R}^{2}$. Assume that the true distribution $F$ has compact support $S\subset\mathbb{R}^{2}$. Assume the following conditions hold:
@@
  • The perturbation size $\delta_{m}$ satisfies $\delta_{m}=o((\log m)^{-c})$, and $\sup_{1\leq i\leq n}\|\hat{Z}_{i}-Z_{i}\|\leq\delta_{m}$ with probability tending to 1 as $m\to\infty$, for some constant $c>0$.
  • For each $k=1,2$, the bandwidth $h_{m}^{(k)}$ for $\hat{F}_{k\,|\,1:(k-1),m}$ is chosen as $h_{m}^{(k)}=h_{m}=(\log m)^{-c}$ for the same constant $c>0$.
@@
For notational simplicity within this proof, let $F_{k}(\cdot)$ denote the marginal CDF of $Z_{k}$, and $F_{k\,|\,1:(k-1)}(\cdot\,|\,u)$ denote the true conditional CDF for the $k$-th dimension given $u=(z_{1},\dots,z_{k-1})$. Similarly, $\hat{F}_{k,m}$ and $\hat{F}_{k\,|\,1:(k-1),m}(\cdot\,|\,u)$ are their estimators. Inverse functions are denoted with 1-1. Let $S_{k}$ represent the compact support of $Z_{k}$ (derived from the compact support $S$ of $F$), and $S_{1:(k-1)}$ be the compact support for the conditioning variables $(Z_{1},\dots,Z_{k-1})$. Because $F$ has compact support $S\subset\mathbb{R}^{d}$, all $Z_{ik}$ and $(Z_{i1},\dots,Z_{i,k-1})$ almost surely lie within their compact projected supports. This is crucial for applying uniform convergence results from previous lemmas, which hold over compact domains.
@@
$\tilde{Z}_{i1}=\hat{F}_{1,m}^{-1}(G(\hat{F}_{1,m}(\hat{Z}_{i1})+e_{i}^{(1)})).$
@@
$\tilde{Z}_{i1}\xrightarrow{d}Y_{1}\sim F_{1}\quad\text{as }m\to\infty.$
Step 2 – convergence of the second coordinate $\tilde{Z}_{i2}$ (conditioning on $\tilde{Z}_{i1}$):
We now analyze the conditional distribution of $\tilde{Z}_{i2}$ given $\tilde{Z}_{i1}=u_{1}$, for any fixed $u_{1}\in S_{1}$ (the compact support of $Y_{1}$). The quantity of interest is $\mathbb{P}(\tilde{Z}_{i2}\leq x_{2}\,|\,\tilde{Z}_{i1}=u_{1})$. For this analysis, we treat $u_{1}$ as a deterministic conditioning variable. The definition of $\tilde{Z}_{i2}$ involves conditioning on $\tilde{Z}_{i1}$, so we consider the variable:
$\tilde{Z}_{i2}(u_{1})=\hat{F}_{2\,|\,1,m}^{-1}(G(\hat{F}_{2\,|\,1,m}(\hat{Z}_{i2}\,|\,\hat{Z}_{i1})+e_{i}^{(2)})\,|\,u_{1}).$
@@
$\bar{Z}_{i2}(u_{1})=\hat{F}_{2\,|\,1,m}^{-1}(G(\hat{F}_{2\,|\,1,m}(Z_{i2}\,|\,Z_{i1})+e_{i}^{(2)})\,|\,u_{1}),$
$T_{i2}(u_{1})=F_{2\,|\,1}^{-1}(G(F_{2\,|\,1}(Z_{i2}\,|\,Z_{i1})+e_{i}^{(2)})\,|\,u_{1}).$
Our goal is to show $\tilde{Z}_{i2}(u_{1})-\bar{Z}_{i2}(u_{1})\xrightarrow{\mathbb{P}}0$ and $\bar{Z}_{i2}(u_{1})-T_{i2}(u_{1})\xrightarrow{\mathbb{P}}0$.
To show $\bar{Z}_{i2}(u_{1})-T_{i2}(u_{1})\xrightarrow{\mathbb{P}}0$, we directly apply the logic from Lemma B.4 to the conditional setting. For a fixed $u_{1}$, the functions $\hat{F}_{2\,|\,1,m}(\cdot\,|\,u_{1})$ and $F_{2\,|\,1}(\cdot\,|\,u_{1})$ are well-defined. Let $q_{C}=G(\hat{F}_{2\,|\,1,m}(Z_{i2}\,|\,Z_{i1})+e_{i}^{(2)})$ and $q_{D}=G(F_{2\,|\,1}(Z_{i2}\,|\,Z_{i1})+e_{i}^{(2)})$. We want to show $\big|\hat{F}_{2\,|\,1,m}^{-1}(q_{C}\,|\,u_{1})-F_{2\,|\,1}^{-1}(q_{D}\,|\,u_{1})\big|\xrightarrow{\mathbb{P}}0$. This splits into:
$\big|\hat{F}_{2\,|\,1,m}^{-1}(q_{C}\,|\,u_{1})-F_{2\,|\,1}^{-1}(q_{C}\,|\,u_{1})\big|+\big|F_{2\,|\,1}^{-1}(q_{C}\,|\,u_{1})-F_{2\,|\,1}^{-1}(q_{D}\,|\,u_{1})\big|.$
  • Term 1: By Theorem B.2, $\sup_{u\in S_{1},q\in[a,b]}\big|\hat{F}_{2\,|\,1,m}^{-1}(q\,|\,u)-F_{2\,|\,1}^{-1}(q\,|\,u)\big|\xrightarrow{\mathbb{P}}0$. Since $u_{1}$ is fixed in $S_{1}$ and $q_{C}$ is a random variable in $[0,1]$, this term goes to 0 in probability, using the same type of proof as in Theorem B.2.
  • Term 2: By uniform continuity of $F_{2\,|\,1}^{-1}(\cdot\,|\,u_{1})$ w.r.t. its first argument, this term goes to 0 in probability if $|q_{C}-q_{D}|\xrightarrow{\mathbb{P}}0$. $|q_{C}-q_{D}|=\big|G(\hat{F}_{2\,|\,1,m}(Z_{i2}\,|\,Z_{i1})+e_{i}^{(2)})-G(F_{2\,|\,1}(Z_{i2}\,|\,Z_{i1})+e_{i}^{(2)})\big|$. By uniform continuity of $G$, this goes to 0 if $\big|\hat{F}_{2\,|\,1,m}(Z_{i2}\,|\,Z_{i1})-F_{2\,|\,1}(Z_{i2}\,|\,Z_{i1})\big|\xrightarrow{\mathbb{P}}0$. This is true by Lemma B.2.
Combining the previous two results, we have $\bar{Z}_{i2}(u_{1})-T_{i2}(u_{1})\xrightarrow{\mathbb{P}}0$.
On the other hand, the term $\tilde{Z}_{i2}(u_{1})-\bar{Z}_{i2}(u_{1})\xrightarrow{\mathbb{P}}0$ involves changes in the inner argument’s conditioning variable ($\hat{Z}_{i1}$ to $Z_{i1}$) and the value being fed into $G$ ($\hat{Z}_{i2}$ to $Z_{i2}$).
$|\tilde{Z}_{i2}(u_{1})-\bar{Z}_{i2}(u_{1})|=\big|\hat{F}_{2\,|\,1,m}^{-1}(G(\hat{F}_{2\,|\,1,m}(\hat{Z}_{i2}\,|\,\hat{Z}_{i1})+e_{i}^{(2)})\,|\,u_{1})-\hat{F}_{2\,|\,1,m}^{-1}(G(\hat{F}_{2\,|\,1,m}(Z_{i2}\,|\,Z_{i1})+e_{i}^{(2)})\,|\,u_{1})\big|.$
@@
$\big|\hat{F}_{2\,|\,1,m}^{-1}(q_{A}\,|\,u_{1})-\hat{F}_{2\,|\,1,m}^{-1}(q_{B}\,|\,u_{1})\big|.$
By the uniform equicontinuity of $\hat{F}_{2\,|\,1,m}^{-1}(\cdot\,|\,u_{1})$ with respect to its first argument from Theorem B.2, this term will converge to 0 in probability if $|q_{A}-q_{B}|\xrightarrow{\mathbb{P}}0$.
@@
$|q_{A}-q_{B}|=\big|G(\hat{F}_{2\,|\,1,m}(\hat{Z}_{i2}\,|\,\hat{Z}_{i1})+e_{i}^{(2)})-G(\hat{F}_{2\,|\,1,m}(Z_{i2}\,|\,Z_{i1})+e_{i}^{(2)})\big|$
By uniform continuity of $G$, it is sufficient to show $\big|\hat{F}_{2\,|\,1,m}(\hat{Z}_{i2}\,|\,\hat{Z}_{i1})-\hat{F}_{2\,|\,1,m}(Z_{i2}\,|\,Z_{i1})\big|\xrightarrow{\mathbb{P}}0$. By plugging in an intermediate term and calling the triangle inequality, we have
$\big|\hat{F}_{2\,|\,1,m}(\hat{Z}_{i2}\,|\,\hat{Z}_{i1})-\hat{F}_{2\,|\,1,m}(Z_{i2}\,|\,Z_{i1})\big|$ $\leq\big|\hat{F}_{2\,|\,1,m}(\hat{Z}_{i2}\,|\,\hat{Z}_{i1})-\hat{F}_{2\,|\,1,m}(Z_{i2}\,|\,\hat{Z}_{i1})\big|$
$\quad+\big|\hat{F}_{2\,|\,1,m}(Z_{i2}\,|\,\hat{Z}_{i1})-\hat{F}_{2\,|\,1,m}(Z_{i2}\,|\,Z_{i1})\big|.$